OpenStudy (anonymous):

this should be my last question integral of [1/(y lny)]dy

OpenStudy (nikvist):

\[\int\frac{dy}{y\ln{y}}=\int\frac{d(\ln{y})}{\ln{y}}=\ln{\ln{y}}+C\]

OpenStudy (anonymous):

wait, what happened?

OpenStudy (zarkon):

\[\text{Let }u=\ln(y)\] \[du=\frac{1}{y}dy\] \[\int \frac{1}{y\ln(y)}dy=\int\frac{1}{u}du=\ln|u|+c=\ln|\ln(y)|+c\]